Formula of finding no of optical and geometrical isomer

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
Compound has n-chiral centres with no symmetry in the molecule 
Then, No. of Optical Isomers (X)= 2^n 
and, No. of Meso-structures (Y)=0 
Hence total no. of stereo isomers= X+Y= 2^n 
this one is the only one , and there is no such formula to calculate the no. of isomers in organic . 
HOPE IT CLEARS 
ALL THE BEST ,..